# GRPO-MA: Multi-Answer Generation in GRPO for Stable and Efficient Chain-of-Thought Training

### v2

@@ A.4.3 Generation Configure
A.5 More Case Study and Visualization
To provide a more intuitive and in-depth analysis of our model’s performance, this section presents a series of curated case studies and visualizations. These examples encompass a range of key tasks, including object detection (Fig. 4) and trajectory prediction (Fig. 5 and Fig. 6). Our aim is to leverage these concrete scenarios to delve into the model’s behavior, decision-making logic, and inherent strengths and limitations.
@@ A.5 More Case Study and Visualization
A.6 More Experimental Analysis
In this section, we further present some experimental results, including the accuracy reward curves during training, and an analysis of the richness of reward signal.
A.6.1 Accuracy Reward Curve
@@
A.6.2 Richness of Reward Signals
For tasks with binary (0-1) rewards, such as Code, Math, Affordance Prediction, Demand Prediction and Simulator-based Visual Manipulation, we compute the proportion of samples whose total reward is positive, which we refer to as the NoZeroRate. Formally, it is defined as
@@ A.6.2 Richness of Reward Signals
where $\mathbf{1}\{\cdot\}$ denotes the indicator function, which equals $1$ if the condition inside holds and 0 otherwise. Here, $T$ is the total number of time steps, $t$ indexes a specific time step, $K$ is the number of thoughts, $M$ is the number of answers per thought, and $\mathrm{AccR}^{t}_{i,j}$ denotes the accuracy reward associated with the $j$-th answer under the $i$-th thought at time step $t$. A higher NoZeroRate indicates a lower proportion of advantage collapses (where collapse means all advantage values become zero), and a higher proportion of effective gradient information contribution.
The statistical results are presented in Tab. 6. We observe that T4A4 achieves the second-highest proportion of non-zero accuracy rewards across all tasks, only behind T16A1. On the one hand, this indicates that under the T4A4 setting, the answers generated by each thought are largely different. On the other hand, it suggests that the diversity of generated answers can be substantially improved by generating additional answers per thought, as shown by the comparison between T4A4 and T4A1.
@@
A.7 Usage of LLMs

### v3

@@ E.3 Generation Configure
Appendix F Case Study and Visualization
To provide a more intuitive and in-depth analysis of our model’s performance, this section presents a series of curated case studies and visualizations. These examples encompass a range of key tasks, including object detection ( Fig. 3 and Fig. 4) and trajectory prediction (Fig. 5 and Fig. 6). Our aim is to leverage these concrete scenarios to delve into the model’s behavior, decision-making logic, and inherent strengths and limitations.
@@ Appendix F Case Study and Visualization
Appendix G More Experimental Analysis
In this section, we further present some experimental results, including inconsistency analysis, the accuracy reward curves during training, and an analysis of the richness of reward signal.
G.1 Inconsistency Analysis
We quantify the inconsistency between thoughts and answers during training. For a thought $th_{i}$ with $M$ answers, if $\operatorname{sign}(A(th_{i}))\neq\operatorname{sign}(A(ans_{i,j}))$, we mark it as inconsistent. The inconsistency rate is defined as $\text{InconsistencyRate}=\tfrac{1}{KM}\sum_{i=1}^{K}\sum_{j=1}^{M}\mathbb{I}[A(th_{i})A(ans_{i,j})<0]$, where $\mathbb{I}[\cdot]$ denotes the indicator function, which equals $1$ if the condition inside holds and 0 otherwise.
G.2 Accuracy Reward Curve
@@
G.3 Richness of Reward Signals
For tasks with binary (0–1) rewards, such as Code, Math, Affordance Prediction, Demand Prediction, and Simulator-based Visual Manipulation, we measure the proportion of samples whose total reward is positive, which we refer to as the NoZeroRate. Formally, it is defined as
@@ G.3 Richness of Reward Signals
where $\mathbf{1}\{\cdot\}$ denotes the indicator function, which equals $1$ if the condition holds and 0 otherwise. Here, $T$ is the total number of time steps, $t$ indexes a specific time step, $K$ is the number of thoughts, $M$ is the number of answers per thought, and $\mathrm{AccR}^{t}_{i,j}$ denotes the accuracy reward associated with the $j$-th answer under the $i$-th thought at time step $t$. Intuitively, a higher NoZeroRate indicates a lower proportion of advantage collapse (i.e., all advantage values being zero), and thus a larger fraction of samples contributing effective gradient information.
The statistical results are reported in Tab. 10. We observe that T4A4 achieves the second-highest NoZeroRate across all tasks, only behind T16A1. This result suggests that T4A4 strikes a favorable balance between the number of thoughts and the number of answers per thought, leading to a high probability that at least one sampled answer yields a non-zero reward. In particular, compared with T4A1, T4A4 exhibits a substantially higher NoZeroRate, indicating that generating multiple answers per thought significantly reduces the likelihood of advantage collapse.
This comparison also addresses a potential concern regarding whether sampling multiple answers conditioned on the same thought truly produces diverse solutions, rather than near-duplicate ones. If answers sampled from a single thought were largely identical, then increasing the number of answers per thought would be unlikely to introduce higher NoZeroRate, and configurations such as T4A4 and T4A1 would be expected to exhibit similar NoZeroRates. However, under our high-temperature sampling regime, we observe a clear gap between the NoZeroRates of T4A4 and T4A1. This gap indicates that additional answers sampled from the same thought substantially increase the probability of observing at least one non-zero reward, thereby leading to a higher NoZeroRate. This provides indirect but quantitative evidence that multi-sampling under a fixed thought yields diverse and effective answers within our framework.
@@
G.4 Wall-clock Time
We provide wall-clock time to further demonstrate the improvement in training efficiency achieved by the GRPO-MA algorithm. We pick two multimodal tasks and one text reasoning task as examples. The results are shown in Fig. 13, Fig. 14, and Fig. 15. As shown in the figure, GRPO-MA achieves the peak performance of various baselines in a shorter time.
G.5 Grad-norm and GSS Curve
We provide complete Grad Norm and GSS curves, shown in Fig. 16, Fig. 17 and Fig 18. Smaller fluctuations in the Grad Norm curve and lower GSS values indicate fewer gradient spikes during training, resulting in more stable training. The grad norm curve and GSS curve corresponding to GRPO-MA both exhibit smaller fluctuations and GSS values.
G.6 More Results on Code and Math
To further examine the generality of our variance-reduction mechanism, we additionally evaluate GRPO-MA on two widely used mathematical and coding benchmarks: GSM8K (Cobbe et al., 2021) for math reasoning and HumanEval (Chen et al., 2021a) for code generation. These experiments use the same training configuration as our main results, with models trained on the DAPO dataset. For GSM8K, we report the Pass@1 accuracy obtained by running the checkpoint trained for mathematical reasoning. For HumanEval, we report Pass@1 and Pass@5 to assess both direct correctness and sampling-based performance. These findings reinforce our claim that answer-level multi-sampling provides a simple, general, and effective variance-reduction strategy that benefits both mathematical reasoning and code generation tasks. The consolidated results are presented in Table 11.
Table 11: Additional Results on GSM8K (Math) and HumanEval (Code). TN: Number of thoughts; AN: Number of answers per thought. Bold indicates the best performance among GRPO variants.
GSM8K (Math) HumanEval (Code)
Model TN AN Pass@1 Pass@1 Pass@5
GRPO 4 1 62.54 56.10 82.93
GRPO 8 1 64.67 56.83 82.93
GRPO 16 1 65.13 57.80 85.37
GRPO-MA 4 4 70.58 58.66 87.20
@@ G.6 More Results on Code and Math
Figure 13: Wall-clock Time on Object Detection
Figure 14: Wall-clock Time on Trajectory Prediction
Figure 15: Wall-clock Time on Code
Figure 16: Grad Norm and GSS on Object Detection
Figure 17: Grad Norm and GSS on Trajectory Prediction
Figure 18: Grad Norm and GSS on Code
Appendix H Usage of LLMs